### v1

@@ Appendix E Proof of Proposition 3
$(\sigma^{2}_{*},\rho_{*})=\arg\min_{\sigma^{2},\rho}\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n\end{subarray}}\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\sigma^{2}(1+\rho(|\mathcal{F}|-1),$ (175)
with $|\mathcal{C}|=c$, and then applied the decoder in Proposition 1. Note that (175) is equivalent to (17), which shows that the optimum noise parameters do not change. However, the resulting MSE changes as the decoder further normalizes the direct MSE resulted by the parameters in (175).
Appendix F Additional Details on Section 3
In this section, we go over the rigorous proofs of the geometric interpretation provided for the two-user example in Section 3.
Privacy Constraint:
@@ Appendix F Additional Details on Section 3
$\mathbb{P}(\mathbf{x}_{i}+\mathbf{Z}_{i}\in\mathcal{A}|\mathbf{x}_{j}+\mathbf{Z}_{j}=\mathbf{y}_{j})\leq e^{\epsilon}\mathbb{P}(\mathbf{x}_{i}^{\prime}+\mathbf{Z}_{i}\in\mathcal{A}|\mathbf{x}_{j}+\mathbf{Z}_{j}=\mathbf{y}_{j})+\delta$ (176)
for each $i\neq j$, for any fixed $\mathbf{x}_{j},\mathbf{y}_{j}$ and $\forall\mathbf{x}_{i},\mathbf{x}_{i}^{\prime}$. Let $\mathbf{Y}_{i}=\mathbf{x}_{i}+\mathbf{Z}_{i}$ for $i=1,2$. The first step is to quantify the conditional distribution $\mathbf{Y}_{i}|\mathbf{Y}_{j}=\mathbf{y}_{j}$ for $i\neq j$. WLOG assume that $i=1$.
@@
Define $f(\mathbf{x}_{i})=\mathbf{x}_{1}+\rho\mathsf{I}_{d}(\mathbf{y}-\mathbf{x}_{2})$. Then, define a new random variable $\mathbf{W}=f(\mathbf{x}_{1})+\mathbf{N}$, where $\mathbf{N}\sim\mathcal{N}\left(\mathbf{0}_{d},\tilde{\sigma}^{2}\mathsf{I}_{d}\right)$. Note that for any given values of $\mathbf{y}_{2}$, $\mathbf{x}_{2}$, $\mathbf{W}\sim\mathbf{Y}_{1}|\mathbf{Y}_{2}=\mathbf{y}_{2}$, (statistically equivalent). Now, consider the $(\epsilon,\delta)$-DP constraint in (176), which is equivalent to,
@@
Recall that $\mathbb{E}[\|\mathbf{Z}_{i}\|^{2}]=\sigma^{2}d=|\mathbf{Z}_{i}|^{2}$ from the vector representation explained in Section 3. Therefore, the privacy constraint in (176) simplifies to,
$|\mathbf{Z}_{i}|\sin\theta\geq\sigma_{\epsilon,\delta}\sqrt{d}=\hat{\sigma}_{\epsilon,\delta}$ (184)

### v2

@@ Appendix E Proof of Proposition 2
$(\sigma^{2}_{*},\rho_{*})=\arg\min_{\sigma^{2},\rho}\max_{\begin{subarray}{c}\mathcal{U}\subseteq\mathcal{U}_{all}\\ t\leq|\mathcal{U}|\leq n\end{subarray}}\sup_{\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}\in\mathbb{B}^{d}}\sigma^{2}(1+\rho(|\mathcal{U}|-1),$ (200)
while satisfying (3), and then applied the decoder in Proposition 1. Note that (200) is equivalent to (2), which shows that the optimum noise parameters do not change in the biased and unbiased cases. However, the resulting MSE changes as the decoder further normalizes the direct MSE resulted by the parameters in (200).
Appendix F Additional Details on Section III
In this section, we go over the rigorous proofs of the geometric interpretation provided for the two-user example in Section III.
F-A Privacy Constraint
@@
$\mathbb{P}(\mathbf{x}_{i}+\mathbf{Z}_{i}\in\mathcal{A}|\mathbf{x}_{j}+\mathbf{Z}_{j}=\mathbf{y}_{j})$
$\quad\leq e^{\epsilon}\mathbb{P}(\mathbf{x}_{i}^{\prime}+\mathbf{Z}_{i}\in\mathcal{A}|\mathbf{x}_{j}+\mathbf{Z}_{j}=\mathbf{y}_{j})+\delta$ (201)
for each $i\neq j$, $\forall\mathbf{x}_{i},\mathbf{x}_{i}^{\prime}\in\mathbb{B}^{d}$, and $\forall\mathcal{A}\subset\mathbb{R}^{d}$ in the Borel sigma algebra. Let $\mathbf{Y}_{i}=\mathbf{x}_{i}+\mathbf{Z}_{i}$ for $i=1,2$. The first step is to quantify the conditional distribution $\mathbf{Y}_{i}|\mathbf{Y}_{j}=\mathbf{y}_{j}$ for $i\neq j$. WLOG assume that $i=1$.
@@
Define $f(\mathbf{x}_{i})=\mathbf{x}_{1}+\rho\mathsf{I}_{d}(\mathbf{y}-\mathbf{x}_{2})$. Then, define a new random variable $\mathbf{W}=f(\mathbf{x}_{1})+\mathbf{N}$, where $\mathbf{N}\sim\mathcal{N}\left(\mathbf{0}_{d},\tilde{\sigma}^{2}\mathsf{I}_{d}\right)$. Note that for any given values of $\mathbf{y}_{2}$, $\mathbf{x}_{2}$, $\mathbf{W}\sim\mathbf{Y}_{1}|\mathbf{Y}_{2}=\mathbf{y}_{2}$, (statistically equivalent). Now, consider the $(\epsilon,\delta)$-DP constraint in (F-A), which is equivalent to,
@@
Recall that $\mathbb{E}[\|\mathbf{Z}_{i}\|^{2}]=\sigma^{2}d=\|\mathbf{Z}_{i}\|_{\mathcal{H}}^{2}$ from the vector representation explained in Section III. Therefore, the privacy constraint in (F-A) simplifies to,
$\|\mathbf{Z}_{i}\|_{\mathcal{H}}\sin\theta\geq\sigma_{\epsilon,\delta}\sqrt{d}=\gamma_{\epsilon,\delta}$ (209)